Known Matrix:


"A=\begin{bmatrix}
 2 & 1 & 2 \\
 1 & 2 & 2 \\
 1 & 1 & 3
\end{bmatrix}""B=\begin{bmatrix}
 3 & 0 & 2 \\
 0 & 1 & a \\
 0 & 2 & 2a
\end{bmatrix}"


  1. Determine the characteristic equation det(A − λI) = 0 of the matrix above
  2. Determine the eigenvalues of the matrix and the basis of the eigenspace

Note: In matrix B, let the value of a be so that the eigenvalues and the basis of the eigenspace are dependent on a.


1
Expert's answer
2021-11-30T11:19:06-0500

a)


"A=\begin{bmatrix}
 2& 1 & 2 \\
 1 & 2 & 2 \\
1 & 1 & 3
\end{bmatrix}"

"A-\lambda I=\begin{bmatrix}
 2 -\lambda & 1 & 2 \\
 1 & 2-\lambda & 2 \\
1 & 1 & 3-\lambda
\end{bmatrix}"

The characteristic equation


"\begin{vmatrix}
 2 -\lambda & 1 & 2 \\
 1 & 2-\lambda & 2 \\
1 & 1 & 3-\lambda
\end{vmatrix}=0"

"(2 -\lambda)\begin{vmatrix}
 2 -\lambda & 2 \\
 1 & 3 -\lambda
\end{vmatrix}-(1)\begin{vmatrix}
 1 & 2 \\
 1 & 3 -\lambda
\end{vmatrix}"

"+(2)\begin{vmatrix}
 1 & 2-\lambda \\
 1 & 1
\end{vmatrix}"

"=(2-\lambda)((2-\lambda)(3-\lambda)-2)-(3-\lambda-2)"

"+2(1-(2-\lambda))"

"=(2-\lambda)(4-5\lambda+\lambda^2)-1+\lambda+2\lambda-2"

"=8-10\lambda+2\lambda^2-4\lambda+5\lambda^2-\lambda^3+3\lambda-3"

"=-\lambda^3+7\lambda^2-11\lambda+5=0"

"-\lambda^2(\lambda-1)+6\lambda(\lambda-1)-5(\lambda-1)=0"

"-(\lambda-1)(\lambda^2-6\lambda+5)=0"

Matrix "A." The characteristic equation is


"-(\lambda-1)(\lambda-1)(\lambda-5)=0"


"B=\begin{bmatrix}
 3& 0 & 2 \\
 0 & 1 & a \\
0 & 2 & 2a
\end{bmatrix}"

"B-\lambda I=\begin{bmatrix}
 3 -\lambda & 0 & 2 \\
 0 & 1-\lambda & a \\
0 & 2 & 2a-\lambda
\end{bmatrix}"

The characteristic equation


"\begin{vmatrix}
 3 -\lambda & 0 & 2 \\
 0 & 1-\lambda & a \\
0 & 2 & 2a-\lambda
\end{vmatrix}=0"

"(3 -\lambda)\begin{vmatrix}
 1 -\lambda & a \\
 2 & 2a -\lambda
\end{vmatrix}-(0)\begin{vmatrix}
 0 & 2 \\
 2 & 2a -\lambda
\end{vmatrix}"

"+(0)\begin{vmatrix}
 0 & 2 \\
 1-\lambda & a
\end{vmatrix}"

"=(3-\lambda)((1-\lambda)(2a-\lambda)-2a)"

"+2(1-(2-\lambda))"

"=(3-\lambda)(2a-\lambda-2a\lambda+\lambda^2-2a)"

"=(3-\lambda)\lambda(\lambda-(1+2a))=0"

Matrix "B." The characteristic equation is


"(3-\lambda)\lambda(\lambda-(1+2a))=0"

2.

Matrix "A." The characteristic equation is


"-(\lambda-1)(\lambda-1)(\lambda-5)=0"

The eigen values are "\lambda_1=5, \lambda_2=1, \lambda_3=1."

"\lambda_1=5"

"\begin{bmatrix}
 2 -\lambda & 1 & 2 \\
 1 & 2-\lambda & 2 \\
1 & 1 & 3-\lambda
\end{bmatrix}=\begin{bmatrix}
 -3 & 1 & 2 \\
 1 & -3 & 2 \\
1 & 1 & -2
\end{bmatrix}"

"R_1=R_1\/(-3)"


"\begin{bmatrix}
 1 & -1/3 & -2/3 \\
 1 & -3 & 2 \\
1 & 1 & -2
\end{bmatrix}"

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & -1/3 & -2/3 \\
 0 & -8/3 & 8/3 \\
1 & 1 & -2
\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & -1/3 & -2/3 \\
 0 & -8/3 & 8/3 \\
0 & 4/3 & -4/3
\end{bmatrix}"

"R_2=-3R_2\/8"

"\begin{bmatrix}
 1 & -1/3 & -2/3 \\
 0 & 1 & -1 \\
0 & 4/3 & -4/3
\end{bmatrix}"

"R_1=R_1+R_2\/3"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -1 \\
0 & 4/3 & -4/3
\end{bmatrix}"

"R_3=R_3-4R_2\/3"


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -1 \\
0 & 0 & 0
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 1 & 0 & -1 \\
 0 & 1 & -1 \\
0 & 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\
x_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
0 \\
0
\end{bmatrix}"

If we take "x_3=t," then "x_1=t, x_2=t, x_3=t."

The eigen vector is "\begin{bmatrix}
 1 \\
1\\
1
\end{bmatrix}."


"\lambda_2=1"

"\begin{bmatrix}
 2 -\lambda & 1 & 2 \\
 1 & 2-\lambda & 2 \\
1 & 1 & 3-\lambda
\end{bmatrix}=\begin{bmatrix}
 1 & 1 & 2 \\
 1 & 1 & 2 \\
1 & 1 & 2
\end{bmatrix}"

"R_2=R_2-R_1"


"\begin{bmatrix}
 1 & 1 & 2 \\
 0 & 0 & 0 \\
1 & 1 & 2
\end{bmatrix}"

"R_3=R_3-R_1"


"\begin{bmatrix}
 1 & 1 & 2 \\
 0 & 0 & 0 \\
0 & 0 & 0
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 1 & 1 & 2 \\
 0 & 0 & 0 \\
0 & 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\
x_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
0 \\
0
\end{bmatrix}"

If we take "x_2=t,x_3=s," then "x_1=-t-2s, x_2=t, x_3=s."

The eigen vectors are "\begin{bmatrix}
 -1 \\
1\\
0
\end{bmatrix}, \begin{bmatrix}
 -2 \\
0\\
1
\end{bmatrix}"



Matrix "B." The characteristic equation is


"(3-\lambda)\lambda(\lambda-(1+2a))=0"

The eigen values are "\lambda_1=3, \lambda_2=1+2a, \lambda_3=0."

"\lambda_1=3"


"\begin{bmatrix}
 3 -\lambda & 0 & 2 \\
 0 & 1-\lambda & a \\
0 & 2 & 2a-\lambda
\end{bmatrix}=\begin{bmatrix}
 0 & 0 & 2 \\
 0 & -2 & a \\
0 & 2 & 2a-3
\end{bmatrix}"


Swap the rows 1and 2


"\begin{bmatrix}
 0 & -2 & a \\
 0 & 0 & 2 \\
0 & 2 & 2a-3
\end{bmatrix}"


"R_1=R_1\/(-2)"


"\begin{bmatrix}
 0 & 1 & -a/2 \\
 0 & 0 & 2 \\
0 & 2 & 2a-3
\end{bmatrix}"

"R_3=R_3-2R_1"


"\begin{bmatrix}
 0 & 1 & -a/2 \\
 0 & 0 & 2 \\
0 & 0& 3a-3
\end{bmatrix}"

"R_2=R_2\/2"


"\begin{bmatrix}
 0 & 1 & -a/2 \\
 0 & 0 & 1 \\
0 & 0& 3a-3
\end{bmatrix}"

"R_1=R_1+aR_2\/2"


"\begin{bmatrix}
 0 & 1 & 0 \\
 0 & 0 & 1 \\
0 & 0& 3a-3
\end{bmatrix}"

"R_3=R_3-3(a-1)R_2"


"\begin{bmatrix}
 0 & 1 & 0 \\
 0 & 0 & 1 \\
0 & 0& 0
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 0 & 1 & 0 \\
 0 & 0 & 1 \\
0 & 0 & 0
\end{bmatrix}\begin{bmatrix}
 x_1 \\
 x_2 \\
x_3
\end{bmatrix}=\begin{bmatrix}
 0 \\
0 \\
0
\end{bmatrix}"

If we take "x_1=t," then "x_1=t, x_2=0, x_3=0."

The eigen vector is "\begin{bmatrix}
 1 \\
0\\
0
\end{bmatrix}."


"\lambda_2=2a+1"

"\begin{bmatrix}
 3 -\lambda & 0 & 2 \\
 0 & 1-\lambda & a \\
0 & 2 & 2a-\lambda
\end{bmatrix}=\begin{bmatrix}
 2-2a & 0 & 2 \\
 0 & -2a & a \\
0 & 2 & -1
\end{bmatrix}"

"R_1=R_1/(2-2a), a\not=1"

"\begin{bmatrix}
 1 & 0 & 1/(1-a) \\
 0 & -2a & a \\
0 & 2 & -1
\end{bmatrix}"

"R_2=R_2\/(-2a)"

"\begin{bmatrix}
 1 & 0 & 1/(1-a) \\
 0 & 1 & -1/2 \\
0 & 2 & -1
\end{bmatrix}"

"R_3=R_3-2R_2"

"\begin{bmatrix}
 1 & 0 & 1/(1-a) \\
 0 & 1 & -1/2 \\
0 & 0 & 0
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 1 & 0 & 1/(1-a) \\
 0 & 1 & -1/2 \\
0 & 0 & 0
\end{bmatrix}=\begin{bmatrix}
 0 \\
0 \\
0
\end{bmatrix}"

If we take "x_3=t," then "x_1=1\/(a-1), x_2=1\/2, x_3=1."

The eigen vector is "\begin{bmatrix}
 1 /(a-1)\\
1/2\\
1
\end{bmatrix}."


"\lambda_3=0"

"\begin{bmatrix}
 3 -\lambda & 0 & 2 \\
 0 & 1-\lambda & a \\
0 & 2 & 2a-\lambda
\end{bmatrix}=\begin{bmatrix}
 3 & 0 & 2 \\
 0 & 1 & a \\
0 & 2 & 2a
\end{bmatrix}"

"R_1=R_1\/3"

"\begin{bmatrix}
 1 & 0 & 2/3 \\
 0 & 1 & a \\
0 & 2 & 2a
\end{bmatrix}"

"R_3=R_3-2R_2"


"\begin{bmatrix}
 1 & 0 & 2/3 \\
 0 & 1 & a \\
0 & 0 & 0
\end{bmatrix}"

Solve the matrix equation


"\begin{bmatrix}
 1 & 0 & 2/3 \\
 0 & 1 & a \\
0 & 0 & 0
\end{bmatrix}=\begin{bmatrix}
 0 \\
0 \\
0
\end{bmatrix}"

If we take "x_3=t," then "x_1=-2\/3, x_2=-a, x_3=1."

The eigen vector is "\begin{bmatrix}
 -2/3\\
-a\\
1
\end{bmatrix}."